OpenStudy (anonymous):

calculate the mass of 250 ml of carbon dioxide that has a density of 2.095 g/L. what would the density of the gas be if the same sample were in a 150 mL volume?

OpenStudy (fretje):

density I name Rho Rho = 2.095 g/L Volume is 0.250L Density equals Mass over Volume Rho = m/V => m = Rho*V = 2.095g/L*0.25L = 0.52375 gram

OpenStudy (fretje):

New density I name Rho2, the new volume V2 Rho2 = m/V2 = 0.52375 g/0.15L = 3.491|66666666666 g/L = +- 3.492 g/L

OpenStudy (fretje):

sorry, I didn't do error definition, so it is an approximation.